# FedFDP: Federated Learning with Fairness and Differential Privacy

### v1

@@ A.2.1 Privacy Loss of $\mathbf{w}_{t+1}^{i}$
(RDP privacy budget of SGMMironov et al. [2019]). Let $SG_{q,\sigma}$, be the Sampled Gaussian Mechanism for some function $f$. If $f$ has sensitivity 1, $SG_{q,\sigma}$ satisfies $(\alpha,R)$-RDP whenever
$R\leq\frac{1}{\alpha-1}\log\max(A_{\alpha}(q,\sigma),B_{\alpha}(q,\sigma))$ (23)
@@
$\left\{\begin{array}[]{l}A_{\alpha}(q,\sigma)\triangleq\mathbb{E}_{z\sim\vartheta_{0}}\left[\left(\vartheta(z)/\vartheta_{0}(z)\right)^{\alpha}\right]\\ B_{\alpha}(q,\sigma)\triangleq\mathbb{E}_{z\sim\vartheta}\left[\left(\vartheta_{0}(z)/\vartheta(z)\right)^{\alpha}\right]\end{array}\right.$ (24)
@@
Finally, the existing work Mironov et al. [2019] describes a procedure to compute $A_{\alpha}(q,\sigma)$ depending on integer $\alpha$.
$A_{\alpha}=\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-k}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma^{2}}\right)$ (25)
@@
(Composition of RDPMironov [2017]). For two randomized mechanisms $f,g$ such that $f$ is $(\alpha,R_{1})$-RDP and $g$ is $(\alpha,R_{2})$-RDP the composition of $f$ and $g$ which is defined as $(X,Y)$(a sequence of results), where $X\sim f$ and $Y\sim g$, satisfies $(\alpha,R_{1}+R_{2})-RDP$
@@
A.2.2 Privacy Loss of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$
Theorem 4.
$R_{loss}^{i}(\alpha)=\frac{T}{\alpha-1}\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-i}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma_{l}^{2}}\right),$ (26)
The proof is similar to that of the $w_{t+1}^{i}$, so we omit it.
A.2.3 Proof Loss of FedFDP
Since both $w_{T}^{i}$ and $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ access the training set, we need to combine their RDP sequentially using Definition 3, and then use Lemma 2 to convert it to $(\epsilon,\delta)$-DP. Therefore, the Theorem 2 is hold.
Lemma 2.
(Conversion from RDP to DP Balle et al. [2020]). if a randomized mechanism $f:D\rightarrow\mathbb{R}$ satisfies $(\alpha,R)$-RDP ,then it satisfies$(R+\ln((\alpha-1)/\alpha)-(\ln\delta+\ln\alpha)/(\alpha-1),\delta)$-DP for any $10<\delta<$ 1.
A.3 Experiment Settings
A.3.1 Parameters Settings
For all datasets, we adopted a widely used heterogeneous settingsLin et al. [2020]Li et al. [2021a] to 10 clients which controlled by a Dirichlet distribution denoted as $Dir(\beta)$, where a smaller $\beta$ indicates a higher degree of heterogeneity. We set the default value of $\beta=0.1$ Lin et al. [2020]Wang et al. [2020].
In all datasets, we use a 4-layer CNN architecture McMahan et al. [2017] which consists of two convolutional layers and two fully connected layers. For FedFair and its baseline algorithms, we set learning rate $\eta=0.1$. For FedFDP and its baseline algorithms, we set the $\eta=1.0$, the batch sample ratio $q=0.05$, clipping bound for gradient $C=0.1$, the noise multipliers for gradient $\sigma=2.0$, the privacy budget $\epsilon=3.52$ and the $\delta=1e-5$. In particular, for the loss values in the FedFDP algorithm that additionally require differential privacy processing, we set the clipping bound for loss $C_{l}=2.5$ and the noise multipliers for loss $\sigma_{l}=5.0$.
In all experiments, we use 25% of the data will form the test dataset, and the remaining 75% of the data will be used for training. All tasks will be repeated five times using different random seeds, and the mean and standard deviation will be reported.
A.3.2 Baselines introduction
The baseline methods included in the comparison are:
  • FedAvgMcMahan et al. [2017]: FedAvg is the most classic algorithm in federated learning. In fact, we can set $\lambda$ to 0, transforming FedFDP into the FedAvg algorithm.
  • SCAFFLODKarimireddy et al. [2020]: SCAFFOLD uses control variates (variance reduction) to correct for the ’client-drift’ in its local updates. It requires significantly fewer communication rounds and is not affected by data heterogeneity or client sampling.
  • FedProxLi et al. [2020a]: FedProx introduces a proximal term to the objective function to address the issues of systems heterogeneity and statistical heterogeneity in federated learning.
  • FedDynAcar et al. [2021]: FedDyn adds linear and quadratic penalty terms to the objective function to solve the problem of inconsistency between the minima of the local-device level empirical loss and those of the global empirical loss.
  • ALI-DPFLLing et al. [2023]: ALI-DPFL dynamically obtains the optimal number of epochs for each round of communication by minimizing the convergence upper bound, thereby addressing the imbalance between local privacy budgets and communication resources.
  • q-FFLLi et al. [2020b]: q-FFL amplifies the loss values of clients, equivalent to giving larger update weights to poorly performing clients, in order to cater to the underperforming clients, uniforming model performance across multiple devices.
A.3.3 Herterogeneity and Scalability
As shown in Table 4 and 5, we conducted additional experiments on the MNIST, FashionMNIST, and CIFAR10 datasets to compare the performance of FedFDP with that of six other algorithms in terms of Heterogeneity and Scalability. The Heterogeneity experiments were mainly conducted in three heterogeneous scenarios, extending from the default $Dir(0.1)$ to include $Dir(0.5)$ and $Dir(1)$. Scalability was expanded from the default 10 clients to include 20 and 50 clients.
Heterogeneities. Tables 4 and 5 show that when confronted with diverse data distributions exhibiting varying levels of heterogeneity, both FedFair and FedFDP outperform baseline methods in accuracy while maintaining significant fairness. With the exception of the $Dir(0.5)$ scenario, where they fall slightly short of q-FFL, they respectively demonstrated fairness improvements of 17.0% and 50% for FedFair and 30.3% and 18.9% for FedFDP compared to the best baselines in $Dir(0.1)$ and $Dir(1.0)$.
Scalability. Table 4 demonstrates that, with accuracy mostly superior to baselines, FedFair exhibited fairness improvements of 9.5%, 15.4%, and 0.01% across three client counts. Table 5 indicates that as the number of clients increases, FedFDP outperformed the best baselines by 0.36%, 0.18%, and 0.59% in accuracy, except in the 50 clients scenario, where it demonstrated fairness improvements of 27.8% and 24.1%.

### v2

@@ A.2. Proof of Theorem 1
(RDP privacy budget of SGM(Mironov et al., 2019)). Let $SG_{q,\sigma}$, be the Sampled Gaussian Mechanism for some function $f$. If $f$ has sensitivity 1, $SG_{q,\sigma}$ satisfies $(\alpha,R)$-RDP whenever
(31) $R\leq\frac{1}{\alpha-1}\log\max(A_{\alpha}(q,\sigma),B_{\alpha}(q,\sigma))$
@@
(32) $\left\{\begin{array}[]{l}A_{\alpha}(q,\sigma)\triangleq\mathbb{E}_{z\sim\vartheta_{0}}\left[\left(\vartheta(z)/\vartheta_{0}(z)\right)^{\alpha}\right]\\ B_{\alpha}(q,\sigma)\triangleq\mathbb{E}_{z\sim\vartheta}\left[\left(\vartheta_{0}(z)/\vartheta(z)\right)^{\alpha}\right]\end{array}\right.$
@@
Finally, the existing work (Mironov et al., 2019) describes a procedure to compute $A_{\alpha}(q,\sigma)$ depending on integer $\alpha$.
(33) $A_{\alpha}=\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-k}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma^{2}}\right)$
@@
(Composition of RDP(Mironov, 2017)). For two randomized mechanisms $f,g$ such that $f$ is $(\alpha,R_{1})$-RDP and $g$ is $(\alpha,R_{2})$-RDP the composition of $f$ and $g$ which is defined as $(X,Y)$(a sequence of results), where $X\sim f$ and $Y\sim g$, satisfies $(\alpha,R_{1}+R_{2})-RDP$